Evaluate subject to the functional relation x=t3 and y=t2.

1
Expert's answer
2021-06-11T10:29:04-0400

Given x = t3 and y = t2

Differentiation is a functional relation of x and y.

So, Differentiating x and y with respect to t

dx/dt = 3t2

dy/dt = 2t

Now functional relation is, dy/dx = (dy/dt)/(dx/dt) = 2t/3t2 = 2/3t

Thus, the functional relation of x and y is

dy/dx = 2/3t